[litho] Morning Litho Status 4/9/13
Gary J Sosa
gsosa at stanford.edu
Tue Apr 9 07:47:28 PDT 2013
Good morning All..
-Karl Suss2- system will not power up. Troubleshooting in progress
General Information/Lab Activities
Tuesday, April 09, 2013
ASML Stepper- Monthly PM rescheduled for Wednesday(4/10) and Thursday(4/11)
Mahnaz- Out at 1:00PM
Will not power up
Does not work as standard aligner. Suspect software or configuration
Per vendor, top optics may need to be configured
Sender not resetting
Receiver not moving up
Continuous testing and investigating
1.Chillplate temp control
2. Wafer Handling
1. Replaced water pump. Testing in progress.
2. Testing various paths and wafer sizes.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the litho